need help solving
-16 +6(z-2) = 2(z-8) + 4z -12

Found 2 solutions by Alan3354, tommyt3rd:
Answer by Alan3354(69443)

need help solving
-16 +6(z-2) = 2(z-8) + 4z -12
--------------
Add 16 to each side.
Then eliminate the parentheses.
go from there.

Answer by tommyt3rd(5050)

-16 +6(z-2) = 2(z-8) + 4z -12
-16+6z-12=2z-16+4z-12
-16+16+6z-12+12=2z-16+16+4z-12+12
0=0

This is an identity - it is always true